OpenStudy (anonymous):

A laptop computer is purchased for $1400 . After each year, the resale value decreases by 25% . What will the resale value be after 5 years?

OpenStudy (nurali):

Year 1 (35% of 1400= 490) 1400- 490= 910 Year 2 (35% of 910 = 318.5) 910- 318.5= 591.5 Year 3 (35% of 591.5= 207 * rounded to the nearest hundredth*) 591.5- 207= 384.5 Year 4 (35% of 384.5 = 134.6 *rounded to the nearest hundredth* ) 384.5 - 134.6= 249.9. Year 4 (35% of 249.9 = 87.5 *rounded to the nearest hundredth* ) 249.9 - 87.5= 162.4

TheSmartOne (thesmartone):

Welcome to Open Study!! :)